OpenStudy (anonymous):

Simplify: 5x - 6y + 4x

OpenStudy (anonymous):

Simplifying 5x + -6y + 4x = 0 Reorder the terms: 5x + 4x + -6y = 0 Combine like terms: 5x + 4x = 9x 9x + -6y = 0 Solving 9x + -6y = 0 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '6y' to each side of the equation. 9x + -6y + 6y = 0 + 6y Combine like terms: -6y + 6y = 0 9x + 0 = 0 + 6y 9x = 0 + 6y Remove the zero: 9x = 6y Divide each side by '9'. x = 0.6666666667y Simplifying x = 0.6666666667y